06/Sep/99 Owl Shift Summary

Valeri Jejer (jejer@uvahea.phys.virginia.edu)
Mon, 6 Sep 1999 08:44:01 -0400 (EDT)

Shift Summary for owl shift, 06/Sep/99

On Shift: Jason L., Valeri.

Hours of beam available at all intensities: 0.0
Hours of beam available at nominal intensity: 0.0
Hours of E832 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 0.6

Accounting of lost time:

Accelerator cryo problems continued from the previous shift.

========================================================================
Runs ended this shift:
Run 14434 - - LASER_SCAN

Tapes brought to FCC this shift:
QKC(K,S)396-398, QKB139
(including tapes for online split)

========================================================================
Accesses during shift:

Hogan and Jason L. to fix laser.

========================================================================
Overview of shift:

No beam entire shift. Got DDD3 veto once, laser illumination problems.
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14434 00:00 Accelerator cryo problems continue.
00:35 Long laser scan ended, stop the run. Tapes are 86%
full, so we change them since there is no beam.
00:40 Init and start BEAM_832 run.
01:16 Ch.13 promises beam in 45 min.
01:30 MCR calls - new estimate 1 hour.
02:20 Jason notices that we are not taking triggers -
DDD3 veto is on for no reason. We decided to stop
run.
02:25 MCR calles - the cryo problem is more serious than
they
thought, no beam for at least two hours.
02:35 Run stop dies. Do drt_fresh_start with tape
overwriting
since we did not have any beam.
04:40 MCR calls saying they have to make an access an warm
up the cryo line. Minimum estimate for beam 2 hours.
Ch. 13 gives a 3 hour estimate.
05:10 Jason notices that there are no CsI laser
illumination
plots. Trying resetting pipeline, drt_fresh_start,
kfrends - nothing helps.
05:40 Ed calls, Jason under his directions checks a few
things, finally Ed recommends to call Julie. Call
Julie, she tells us to call Hogan, who is on his
way.
05:48 Cryo line valve has ben replaced, cooling down will
take at least 2 hours.
06:40 Hogan and Jason L. go on access to look at the
laser.
07:10 Hogan and Jason are back. the problem was filter in
wrong position blocking the laser beam.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift